OpenStudy (anonymous):

can any one tell me what's the reason/logic behind 7^0=1 and what is the case with 0^0?

OpenStudy (anonymous):

0^0 is undefined

OpenStudy (anonymous):

It is defined as \[7^{0}=1\] nobody can't change

OpenStudy (anonymous):

1^1 =1 2^1 =1 1^0 =1 2^0= 1 0^0=? 0^0=?

OpenStudy (anonymous):

yes, \[n^0=1\]as long as n is not equal to 0

OpenStudy (anonymous):

0^1=? 0^0=?

OpenStudy (anonymous):

0^1 = 0

OpenStudy (anonymous):

Right

OpenStudy (anonymous):

what's the reason behind 7^0=1 -^0=1 but 0^0=0 how????

OpenStudy (anonymous):

0^0 is not 0

OpenStudy (anonymous):

0 to any power is 0 any power of 0 is 1 so is 0^0 answer 1 or 0 we call it undefined

OpenStudy (anonymous):

7 ^3 / 7^3 = 1 = 7^(3-3) = 7^0 therefore 7^0 = 1 -this is true for any number except 0^0 which is indeterminate

OpenStudy (anonymous):

Any number (except zero) divided by itself equals 1. \[x ^{m}\div x ^{n} = x ^{m - n}\] \[7 ^{m}\div 7 ^{n} = 1\] \[7 ^{n}\div 7 ^{n} = 7 ^{n - n}\]\[1= 7^{0}\]

OpenStudy (anonymous):

xerxes Cool Name

OpenStudy (anonymous):

erratum: \[7^{n}\div 7^{n} = 1\]

OpenStudy (anonymous):

great work xerxes. now explain the deal with 0^0

OpenStudy (anonymous):

\[x ^{n}\div x ^{n} = x ^{0}\] let x = 0... the quotient is 0^0. But the divisor is zero! and division by zero is NOT permissible in math. my teacher said NEVER DIVIDE BY ZERO.

OpenStudy (anonymous):

that makes 0^0 indeterminate.